OpenStudy (anonymous):

why is 1+ tan^2θ = sec^2θ ??

OpenStudy (irishboy123):

because \(\cos^2 \theta + \sin^2 \theta = 1\) because Pythagoreas Theorem holds water....🍀 multiply your equation by \(\cos^2 \theta\) and you will get to the original identity.

OpenStudy (anonymous):

Ohh I see! Thank you @IrishBoy123